# RingID: Rethinking Tree-Ring Watermarking for Enhanced Multi-Key Identification

### v2

@@ 1 Introduction
With the advancement and popularity of diffusion models, a vast number of high-quality diffusion generated images are circulating on the internet. It has become increasingly important to identify AI-generated images to detect fabrication and to track and protect copyrights. Luckily, watermarking [35, 1, 9] offers a reliable solution.
@@ 2.1 Diffusion Models
Diffusion models  [15, 29, 8] signify a swiftly advancing category of generative models. They define a forward diffusion process that gradually add small amount of Gaussian noise to the input data and learn to recover the clean data from the distortion [15]. Diffusion models initially demonstrate huge potential in generating high-resolution images [15, 30], resulting in globally recognized products like Midjourney [22], DALL-E [3, 24] and StableDiffusion [27]. Thereafter, they are broadly applied to modeling various types of data, e.g. video [36, 13, 28, 4, 39], audio [16], 3d object [25, 19], humans [33, 6], etc.
@@ 2.2 Watermarking Diffusion Models
When it comes to watermarking diffusion models, there are two different types of objectives: watermarking the model weights [2, 34, 40, 20, 43] or the generated content [38, 44, 32, 41, 10]. Model weights watermarking methods aim to protect the intellectual property of the model itself and identify whether another model is an instance of plagiarism. While content watermarking methods protect the copyright of generated images [9, 38] or videos [41, 21] by adding imperceptible perturbations onto pixels. They either imprint the watermark into the frequency via DFT [17], DCT [7] and wavelet transformations [37], or directly onto pixels via an encoder network [44, 32, 12]. Recently, Tree-Ring [35] proposes to imprint a tree-ring watermark pattern into the initial diffusion noise. It demonstrates strong robustness to various attacks. In this work, we delve into Tree-Ring, uncovering its source of power in watermark verification and pinpointing its vulnerability in identifying multiple keys. Subsequently, we introduce RingID, a systematic solution fortified with enhanced identification capability.
@@ 4 Delve into Tree-Ring Watermarking
In this section, we will reveal why Tree-Ring [35] is robust to various image transformations, even some of them are not considered by design. We further evaluate Tree-Ring in the multi-key identification task and pinpoint vulnerabilities in its original design.
@@ 4.2 Distribution Shift in Watermarking
Detailed derivation is elaborated in Appendix A Actually, we find any operation leading to a sufficient distribution shift can be considered a good stand-alone watermarking approach. We demonstrate the operation of discarding the imaginary part performs well on its own in Appendix B.
@@ 4.2.1 Effect in Different Tasks and Attacks
Fig. 3 illustrates the effect of distribution shift under different tasks. In Verification, it helps distinguish between watermarked and non-watermarked images. Next, we go one step further and consider the situation when the image undergoes different attacks like JPEG or Resize. If the tree-ring pattern itself can cope with a certain attack, the improvement brought by the distribution shift will be small, because the two distributions are already separable even without the help of shift. On the contrary, if the tree-ring pattern itself cannot deal with a certain attack, the improvement brought by the distribution shift will be significant. In our ablation (Tab. 7), we find distribution shift helps a lot under Rotation and Crop/Scale attacks. This implies that the original design of tree-ring cannot handle these two attacks. When it comes to Identification, things are a little different. Since different watermarks experience the same distribution shift, discarding the imaginary part doesn’t help distinguish them.
@@ 4.4 Why is Tree-Ring Vulnerable to Rotation and Crop/Scale?
4.4.1 Vulnerable to Crop & Scale
Scaling an image by a factor of $a$ corresponds to compressing the frequency spectrum $X\left[u,v\right]$ in both magnitudes and positions: $\frac{1}{a^{2}}X\left[\frac{u}{a},\frac{v}{a}\right]$. Cropping an image modulates the frequency magnitudes by a sinc function. It is straightforward that these transformations easily disrupt the tree-ring pattern on frequency domain. As shown in Fig. 4(e), the pattern is completely disrupted.
4.4.2 Vulnerable to Rotation
Although Tree-Ring is designed as ring patterns to seek rotation robustness, we find the existing design ignores many details, leading to no rotation robustness. In short, a combination of corner cropping when rotating, a lossy injection process, and the rotationally asymmetric shape cause the problem. We will explain in detail and propose solutions in Sec. 5.2.
@@ 5.2 Being Invariant to Rotation
Although the tree-ring pattern is designed to be rotationally symmetric, it exhibits unexpected vulnerability to rotation transformation as shown in Tab. 2. Below, we locate the causes of the problems and propose a series of improvements to solve them.
@@ 5.2.2 Lossless Imprinting
The operation of discarding the imaginary part during watermarking not only shifts the distribution of $\ell_{1}$-to-reference, but also disrupt the pattern, resulting in the lose of its rotational symmetry. Fig. 4(d) visualizes the actually carried pattern. To be more concrete, recall from Fig. 2, when the watermarked noise $X[u,v]$ is transformed back to the spatial $x[m,n]$, only its real part $x_{\text{re}}[m,n]$ is kept for image generation. From a frequency domain perspective, this is equivalent to generation with only the information from its conjugate symmetry $X_{\text{cs}}[u,v]$, which is not necessarily equal to $X[u,v]$. To leverage the rotational symmetry of the ring patterns, we attempt to eliminate the information loss during imprinting by solving $X[u,v]=X_{\text{cs}}[u,v]$. The results indicate that $X_{\text{re}}[u,v]$ should be an even function and $X_{\text{im}}[u,v]$ should be an odd function about the Fourier center. To satisfy these constraints, we imprint the tree-ring pattern only in the real part and leave the imaginary part blank. We also align the center of the tree-ring pattern with the Fourier center $(32,32)$. Originally, the tree-ring pattern was put at $(31,32)$ and imprinted on both the real and the imaginary.
5.2.3 Make A Rounder Ring
@@ 6.1 Experimental Setup
Following Tree-Ring [35], we experiment with the state-of-the-art opensource diffusion model StableDiffusion-V2 [27]. By default, RingID imprints a tree-ring watermark with radius 3-14 on the channel 3 of the initial noise and a Gaussian noise watermark on channels 0. We adopt discretization (to 64 and -64), spatial shift, rounder drawing, and lossless imprinting as ordered in Fig. 2. We report the ROC-AUC between 1000 watermarked and 1000 unwatermarked images as the verification metric. In identification, we report identification accuracy with different numbers of keys. Under both verification and identification settings, the evaluations are performed across various image distortions, including rotation, JPEG compression, cropping and scaling (C&S), blurring, noising, and brightness adjustment. We assess the generation quality with CLIP scores [26] by OpenCLIP-ViT/G [5] between the generated image and the prompts to measure text-image alignment, and Frechet Inception Distance (FID) [14] to quantify the similarity between generated and natural images. we generate 5000 images and evaluate FID on the COCO 2017 training dataset [18]. We set both generation and inversion steps to 50.
@@ 6.2 Comparison with Baselines
6.2.1 Robustness
In verification (Tab. 1), both RingID and Tree-Ring perfectly distinguish between watermarked and unwatermarked images when no attacks are performed, achieving AUC score of 1. However, RingID exhibit stronger robustness to various image distortions than Tree-Ring, achieving near-perfect average AUC score, even though it is not optimized for verification. In the more challenging identification task, (Tab. 2), RingID significantly outperforms Tree-Ring in all settings. When the number of keys to identify is 32, Tree-Ring’s average identification accuracy (excluding C&S) is 0.465, while RingID’s accuracy reaches up to 0.992. As the number of keys increases to 2048, this gap widens drastically (0.077 v.s. 0.942), further highlighting RingID’s superior robustness and larger valid key capacity.
@@ 6.2.1 Robustness
6.2.2 Image Quality
RingID achieves strong performance without compromising the generation quality. It maintains comparable text-image alignments with Tree-Ring. Both methods achieve similar CLIP scores (0.364 v.s. 0.365). RingID obtains an FID of 26.13, which is also comparable to Tree-Ring’s FID of 25.93.
@@ 6.3 Ablation Study
6.3.1 Components of RingID
We study the effectiveness of all proposed components of RingID. From Tab. 3, we can find that all designs are crucial for the final success of RingID. Spatial shift, lossless imprinting, and discretization play a significant role in the robustness of rotation. While the use of rounder rings also increase the accuracy by approximately 40% (from 0.620 to 0.860). Notably, spatial shift sacrifices the robustness to blur and brightness attacks for rotation robustness.
6.3.2 Heterogeneous Watermarking
RingID employs the combination of a Gaussian noise watermark and a ring watermark. Here, we ablate the choice of imprinting the noise watermark on different channels. By comparing the first 3 rows of Tab. 4, we find that heterogeneous watermarking can effectively combine the advantage from both the noise watermark and the ring watermark. Among single-channel configurations, imprinting noise watermark on channel 0 (0.819) achieves the best performance. The combination of channels 0 and 1 yields the highest accuracy among all trials. Interestingly, imprinting noise watermarks on all available channels (0, 1, 2) decrease the accuracy from 0.830 to 0.807.
@@ 6.4 Discussion
6.4.1 Where to Put Rings?
In this section, we will discuss where is the best place to put rings. We could imprint all rings on one channel or distribute them evenly across multiple channels. So what is the better practice? In Tab. 5, we fix the total number of rings to 12 and distribute them to 1-4 channels. we can find that performance drops as we distribute rings to more channels. What’s more, imprinting rings on multiple channels results in a greater likelihood of generating artifacts (shown in Appendix C). So given fixed capacity, it’s better to imprint all rings on a single channel.
When we put all rings on a single channel, what’s the best inner radius and outer radius? From Tab. 6 we can see that 3-14 forms a good choice for 11 rings with capacity 2048. The optimal ring radius also varies for different ring numbers.
6.4.2 More Keys
We would like to explore whether RingID can handle a larger number of keys. Fig. 5 shows the average identification accuracy under various attacks v.s. number of keys. We observe that RingID maintains good accuracy even with a large number of keys. And identification accuracy linearly decreases as the number of keys exponentially increase.
@@ 6.5 Qualitative Results
As shown in Fig. 6, both RingID and Tree-Ring watermarking can generate images of good visual quality. Note that unlike low-level watermarking techniques, both RingID and Tree-Ring cause subtle changes in the image layout. However, RingID generates images with layout that is closer to the unwatermarked image. We owe this to avoiding watermarking on the low-frequency part (radius 0-3).
Figure 6: Qualitative results of images with no watermark, Tree-Ring watermark, and RingID watermark. The subtle changes in image layouts show that both methods embed watermarks into the image’s semantic content, such as face orientations.
@@ 7 Conclusion and Limitation
In this work, we conduct a comprehensive investigation on Tree-Ring Watermarking. We reveal that the distribution shift introduced in the watermarking process contributes to the detection of the watermark. Further, we systematically evaluate Tree-Ring Watermarking in the identification task and pinpoint its limitations. Based on the findings and analysis, we propose RingID, which demonstrates enhanced identification capability.
Limitation: Under multiple-key identifications, one limitation of both Tree-Ring and our RingID is the vulnerability to cropping and scaling attacks due to ring pattern disruption. Future work may explore other transform domains to resist these transformations.
Ethics: Like any watermarking methods, RingID could be misused by malicious actors or authoritarian regimes to track, monitor, or persecute individuals based on their creative outputs, which could lead to infringements of privacy.
@@ A.2 Tree-Ring’s Pipeline
As visualized in Figure Fig. 7, Tree-Ring proposes to add a watermark to a frequency domain initial latent noise $X[u,v]$ (of size $N\times N$) by substituting the value of $M$ pixels within a watermark region mask $\mathcal{M}$. For these watermarked pixels $X[u,v]\in\mathcal{M}$ (corresponds with $w$ used in the paper main content), Tree-Ring samples their values from a circularly-symmetric complex normal distribution $\mathcal{N}_{\mathcal{C}}\left(0,N^{2}\right)=\mathcal{N}\left(0,\frac{N^{2}}{2}\right)+j\mathcal{N}\left(0,\frac{N^{2}}{2}\right)$. Other non-watermarked pixels $X[u,v]\notin\mathcal{M}$, individually, also follow this distribution, but the difference is that in Tree-Ring’s context, $X[u,v]\notin\mathcal{M}$ are spatially ensured to be conjugate symmetric $X[u,v]=X_{cs}[u,v]$, while $X[u,v]\in\mathcal{M}$ does not have such a guarantee. During the analysis, we view each pixel as a random variable.
@@ A.5 Distribution Shift in $\ell_{1}$ Distance
Generally speaking, discarding the imaginary part causes the expectation of $\ell_{1}$-to-reference distance to shift by $\frac{\sqrt{3}}{2}$ statistically. This shift factor is derived under several assumptions. In practice, situations are more complicated. Tree-Ring [35] injects a fixed watermark pattern. Thus $\hat{X}$, $\hat{X}^{\prime}$ and $Z$ all carries information of a specific pattern and correlates with each other. They are not i.i.d samples. Both pattern matching and the distribution shift contributes to the expectation of $\ell_{1}$ distance in Eq. 16.
@@ A.6 Distribution Shift in Real Scenarios
In Control 1, we follow the original setup of Tree-Ring [35]. The original setup aims at distinguishing between the recovered watermark $\hat{w}$ (shifted) and the null watermark $\hat{w}_{\varnothing}$ (not shifted) recovered from the unwatermarked images, i.e. $\lVert\hat{w}-w\rVert_{1}$ v.s. $\lVert\hat{w}_{\varnothing}-w\rVert_{1}$.
@@
We compare the results of Control 1 and 2 in Tab. 7. We can find general performance drop under all attacks in Control 2. The average AUC decreases from 0.975 to 0.913, making it more challenging to distinguish watermarked and non-watermarked images without the help of distribution shift. Further observation reveals that the major drop occurs in Rotation and Crop & Scale attacks. This result aligns with our identification results Tab. 2 in the main text. indicating that distribution shift contributes substantially to the robustness to Rotate and Crop & Scale attacks. This also implies that the original tree-ring watermark pattern cannot handle these attacks.

### v3

@@ 1 Introduction
With the popularity of diffusion models, a vast number of high-quality diffusion generated images are circulating on the internet. It has become increasingly important to identify AI-generated images to detect fabrication and to track and protect copyrights. Luckily, watermarking [1, 13] offers a reliable solution.
@@ 2.1 Diffusion Models
Diffusion models  [20, 34, 12] signify a swiftly advancing category of generative models. They define a forward diffusion process that gradually add small amount of Gaussian noise to the input data and learn to recover the clean data from the distortion [20]. Diffusion models initially demonstrate huge potential in generating high-resolution images [20, 35], resulting in globally recognized products like Midjourney [27], DALL-E [3, 29] and StableDiffusion [32]. Thereafter, they are broadly applied to modeling various types of data, e.g. video [42, 17, 33, 4, 47, 36], audio [21], 3d object [30, 24], humans [39, 7, 8, 53], etc.
@@ 2.2 Watermarking Diffusion Models
When it comes to watermarking diffusion models, there are two different types of objectives: watermarking model weights [2, 40, 48, 25, 51] or generated content [46, 52, 38, 49, 14, 6]. Model weights watermarking methods aim to protect the intellectual property of the model itself and identify whether another model is an instance of plagiarism. While content watermarking methods protect the copyright of generated images [13, 46, 45] or videos [49, 26] by adding imperceptible perturbations onto pixels. They either imprint the watermark into the frequency via DFT [22], DCT [9, 11] and wavelet transformations [43, 10], or directly onto pixels via an encoder network [52, 38, 16, 44]. Recently, Tree-Ring [41] proposes to imprint a tree-ring watermark pattern into the initial diffusion noise. It demonstrates strong robustness to various attacks. In this work, we delve into Tree-Ring, uncovering its source of power in watermark verification and pinpointing its vulnerability in identifying multiple keys. Subsequently, we introduce RingID, a systematic solution fortified with enhanced identification capability.
@@ 4 Delve into Tree-Ring Watermarking
In this section, we will reveal why Tree-Ring [41] is robust to various image transformations, even some of them are not considered by design. We further evaluate Tree-Ring in the multi-key identification task and pinpoint its vulnerabilities.
@@ 4.2 Distribution Shift in Watermarking
Detailed derivation is elaborated in Appendix A. Actually, we find any operation leading to a sufficient distribution shift can be considered a good stand-alone watermarking approach. We demonstrate the operation of discarding the imaginary part performs well on its own in Appendix B.
@@ 4.2.1 Effect in Different Tasks and Attacks
Fig. 3 illustrates the effect of distribution shift under different tasks. In Verification, it helps distinguish between watermarked and non-watermarked images. Next, we go one step further and consider the situation when the image undergoes different attacks like JPEG or Resize. If the tree-ring pattern itself can cope with a certain attack, the improvement brought by the distribution shift will be small, because the two distributions are already separable even without the help of shift. On the contrary, if the tree-ring pattern itself cannot deal with a certain attack, the improvement brought by the distribution shift will be significant. Empirically (Sec. A.6), we find distribution shift helps a lot under Rotation and Crop/Scale attacks. This implies that the original design of tree-ring cannot handle these two attacks. When it comes to Identification, things are a little different. Since different watermarks experience the same distribution shift, discarding the imaginary part doesn’t help distinguishing them.
@@ 4.4 Why is Tree-Ring Vulnerable to Rotation and Crop/Scale?
Vulnerable to Crop & Scale Scaling an image by a factor of $a$ corresponds to compressing the frequency spectrum $X\left[u,v\right]$ in both magnitudes and positions: $\frac{1}{a^{2}}X\left[\frac{u}{a},\frac{v}{a}\right]$. Cropping an image modulates the frequency magnitudes by a sinc function. It is straightforward that these transformations easily disrupt the tree-ring pattern on frequency domain. As shown in Fig. 4(e), the pattern is completely disrupted.
Vulnerable to Rotation Although Tree-Ring is designed as ring patterns to seek rotation robustness, we find the existing design ignores many details, leading to no rotation robustness. In short, a combination of corner cropping when rotating, a lossy injection process, and the rotationally asymmetric shape cause the problem. We will explain in detail and propose solutions in Sec. 5.2.
@@ 5.2 Being Invariant to Rotation
Although the tree-ring pattern is designed to be rotationally symmetric, it exhibits unexpected vulnerability to rotation as shown in Tab. 2. Below, we identify the causes of the issue and propose a series of measures to address it.
@@ 5.2.2 Lossless Imprinting
The operation of discarding the imaginary part during watermarking not only shifts the distribution of $\ell_{1}$-to-reference, but also disrupt the pattern, resulting in the loss of its rotational symmetry. Fig. 4(d) visualizes the actually carried pattern. More concretely, recall from Fig. 2, when the watermarked noise $X[u,v]$ is transformed back to the spatial $x[m,n]$, only its real part $x_{\text{re}}[m,n]$ is kept for image generation. From a frequency domain perspective, this is equivalent to generation with its conjugate symmetry $X_{\text{cs}}[u,v]$, which is not necessarily equal to $X[u,v]$. To leverage the rotational symmetry of the ring patterns, we attempt to eliminate the information loss during imprinting by solving $X[u,v]=X_{\text{cs}}[u,v]$. The results indicate that $X_{\text{re}}[u,v]$ should be an even function and $X_{\text{im}}[u,v]$ should be an odd function about the Fourier center. To satisfy these constraints, we imprint the tree-ring pattern only in the real part and leave the imaginary part blank. We also align the center of the tree-ring pattern with the Fourier center $(32,32)$. Originally, the tree-ring pattern was put at $(31,32)$ and imprinted on both the real and the imaginary.
5.2.3 Making A Rounder Ring
@@ 6.1 Experimental Setup
Following Tree-Ring [41], we experiment with the state-of-the-art opensource diffusion model StableDiffusion-V2 [32]. By default, RingID imprints a tree-ring watermark with radius 3-14 on the channel 3 of the initial noise and a Gaussian noise watermark on channels 0. We adopt discretization (to 64 and -64), spatial shift, rounder drawing, and lossless imprinting as ordered in Fig. 2. We report the ROC-AUC between 1000 watermarked and 1000 unwatermarked images as the verification metric. In identification, we report identification accuracy with different numbers of keys. Evaluation is performed under various image distortions, following the original setup of Tree-Ring [41], i.e., JPEG 25, rotation $75^{\circ}$, $75\%$ random C&S, blur kernel 8, noise std 0.1, brightness [0, 6]. We assess the generation quality with CLIP scores [31] by OpenCLIP-ViT/G [5] between the generated image and the prompts to measure text-image alignment, and Frechet Inception Distance (FID) [19] to quantify the similarity between generated and natural images. we generate 5000 images and evaluate FID on the COCO 2017 training dataset [23]. We set both generation and inversion steps to 50.
@@ 6.2 Comparison with Baselines
Robustness In verification (Tab. 1), both RingID and Tree-Ring perfectly distinguish between watermarked and unwatermarked images when no attacks are performed, achieving AUC score of 1. However, RingID exhibit stronger robustness to various image distortions than Tree-Ring, achieving near-perfect average AUC score, even though it is not optimized for verification. In the more challenging identification task, (Tab. 2), RingID significantly outperforms Tree-Ring. When identifying 32 keys, Tree-Ring’s average identification accuracy (excluding C&S) is 0.465, while RingID’s accuracy reaches up to 0.992. As the number of keys increases to 2048, this gap widens drastically (0.077 v.s. 0.942), further highlighting RingID’s superior robustness and larger valid key capacity.
@@
Image Quality RingID achieves strong performance without compromising the generation quality. It maintains comparable text-image alignments with Tree-Ring. Both methods achieve similar CLIP scores (0.364 v.s. 0.365). RingID obtains an FID of 26.13, which is also comparable to Tree-Ring’s FID of 25.93.
@@ 6.3 Ablation Study
Components of RingID Tab. 3 ablates all proposed components of RingID. We can find that all designs are crucial for the final success of RingID. Spatial shift, lossless imprinting, and discretization play a significant role in the robustness of rotation. While the use of rounder rings also increase the accuracy by approximately 40% (from 0.620 to 0.860). Notably, spatial shift sacrifices the robustness to blur and brightness attacks for rotation robustness.
Heterogeneous Watermarking RingID employs the combination of a Gaussian noise watermark and a ring watermark. Here, we ablate the choice of imprinting the noise watermark on different channels. By comparing the first 3 rows of Tab. 4, we find that heterogeneous watermarking can effectively combine the advantage from both the noise watermark and the ring watermark. Among single-channel configurations, imprinting noise watermark on channel 0 (0.819) achieves the best performance. The combination of channels 0 and 1 yields the highest accuracy among all trials. Interestingly, imprinting noise watermarks on all available channels (0, 1, 2) decrease the accuracy from 0.830 to 0.807.
@@ 6.4 Discussion
Where is the Best Place to Put Rings? We could imprint all rings on one channel or distribute them evenly across multiple channels. In Tab. 5, we fix the total number of rings to 12 and distribute them to 1-4 channels. we can find that performance drops as we distribute rings to more channels. What’s more, imprinting rings on multiple channels cause artifacts to appear more frequently (shown in Appendix C). So given fixed capacity, it’s better to imprint all rings on a single channel. This raises another question: what are the optimal inner and outer radii if all rings are placed on a single channel? From Tab. 6 we can see that 3-14 forms a good choice for 11 rings with capacity 2048. The optimal ring radius also varies for different ring numbers.
More Keys We would like to explore whether RingID can handle a larger number of keys. Fig. 5 shows the average identification accuracy under various attacks v.s. number of keys. We observe that RingID maintains good accuracy even with a large number of keys. And identification accuracy linearly decreases as the number of keys exponentially increase.
@@ 6.5 Qualitative Results
As shown in Fig. 6, both RingID and Tree-Ring watermarking can generate images of good visual quality. Unlike low-level watermarking techniques, both RingID and Tree-Ring cause subtle changes in the image layout. However, RingID generates images with layout that is closer to the unwatermarked image. We owe this to avoiding watermarking on the low-frequency part (radius 0-3).
Figure 6: Qualitative results. The subtle changes in image layouts show that both Tree-Ring and RingID embed watermarks into the image’s semantic content.
@@ 7 Conclusion and Limitation
We conduct a comprehensive investigation on Tree-Ring Watermarking, revealing its limitations in identifying multiple keys. We then propose RingID to enhance identification capability.
Limitation: Under multi-key identification, one limitation is that both Tree-Ring and our RingID are vulnerable to cropping and scaling attacks. Future work may explore other transform domains to resist these attacks.
Acknowledgment: This project is supported by the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 (FY2022) Award 22-5406-A0001.
@@ A.2 Tree-Ring’s Pipeline
As visualized in  Fig. 7, Tree-Ring proposes to add a watermark to a frequency domain initial latent noise $X[u,v]$ (of size $N\times N$) by substituting the value of $M$ pixels within a watermark region mask $\mathcal{M}$. For these watermarked pixels $X[u,v]\in\mathcal{M}$ (corresponds with $w$ used in the paper main content), Tree-Ring samples their values from a circularly-symmetric complex normal distribution $\mathcal{N}_{\mathcal{C}}\left(0,N^{2}\right)=\mathcal{N}\left(0,\frac{N^{2}}{2}\right)+j\mathcal{N}\left(0,\frac{N^{2}}{2}\right)$. Other non-watermarked pixels $X[u,v]\notin\mathcal{M}$, individually, also follow this distribution, but the difference is that in Tree-Ring’s context, $X[u,v]\notin\mathcal{M}$ are spatially ensured to be conjugate symmetric $X[u,v]=X_{cs}[u,v]$, while $X[u,v]\in\mathcal{M}$ does not have such a guarantee. During the analysis, we view each pixel as a random variable.
@@ A.5 Distribution Shift in $\ell_{1}$ Distance
Generally speaking, discarding the imaginary part causes the expectation of $\ell_{1}$-to-reference distance to shift by $\frac{\sqrt{3}}{2}$ statistically. This shift factor is derived under several assumptions. In practice, situations are more complicated. Tree-Ring injects a fixed watermark pattern. Thus $\hat{X}$, $\hat{X}^{\prime}$ and $Z$ all carries information of a specific pattern and correlates with each other. They are not i.i.d samples. Both pattern matching and the distribution shift contributes to the expectation of $\ell_{1}$ distance in Eq. 16.
@@ A.6 Distribution Shift in Real Scenarios
In Control 1, we follow the original setup of Tree-Ring. The original setup aims at distinguishing between the recovered watermark $\hat{w}$ (shifted) and the null watermark $\hat{w}_{\varnothing}$ (not shifted) recovered from the unwatermarked images, i.e. $\lVert\hat{w}-w\rVert_{1}$ v.s. $\lVert\hat{w}_{\varnothing}-w\rVert_{1}$.
@@
We compare the results of Control 1 and 2 in Tab. 7. We can find general performance drop under all attacks in Control 2. The average AUC decreases from 0.975 to 0.913, making it more challenging to distinguish watermarked and non-watermarked images without the help of distribution shift. Further observation reveals that the major drop occurs in Rotation and Crop & Scale attacks, indicating that distribution shift contributes substantially to the robustness to Rotate and Crop & Scale attacks. This also implies that the original Tree-Ring watermark pattern cannot handle these attacks.